OpenStudy (anonymous):

how would i change (y-0)=-2(x-2) into the form (y2-y1)=(x2-x1)

OpenStudy (debbieg):

I'm not sure what you mean.... it's already in point-slope form, (y2-y1)=m(x2-x1) ?? You can't change the slope. Are you sure you're reading the problem correctly?

OpenStudy (anonymous):

I don't see y2 or x2 in the original equation either, so really you can't.

OpenStudy (debbieg):

y2 and x2 are just the y and x in the original.

OpenStudy (anonymous):

yea my teacher was giving me tips for it and here is what she said.(y - 0) = -2(x-2) Now can you put it into standard form from here? 2. You can find the slope using (y2-y1)/(x2-x1) and then use point-slope form: y-y1 = m(x-x1) to write the equation.

OpenStudy (debbieg):

OK standard form is: Ax + By = C Where A, B and C are integers. Usually we also want to make sure A>0 and A, B, and C have no common factors (these rules make standard form unique).

OpenStudy (anonymous):

Ah, sorry, haven't done algebra in a while lol

OpenStudy (debbieg):

So far (y-0)=-2(x-2) you want to expand through the ( )'s, and then shuffle terms around as needed to have Ax + By = C

OpenStudy (debbieg):

*for

OpenStudy (debbieg):

It won't take much. just move the x term to the left side so it is with the y, and leave the constant on the right side. :)

OpenStudy (anonymous):

ok so what i need to do is multiply the things in the ( ) by -2 to start?

OpenStudy (debbieg):

On the right side, yes. On the left side just simplify, you don't need "-0" :)

OpenStudy (debbieg):

Tell me what you get after that first step.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i multiplied the right side and got it in y=mx+b form so it is now y=-2x-2

OpenStudy (debbieg):

Double check that distribution on the right side....

OpenStudy (anonymous):

oh i see it needs to be +4

OpenStudy (debbieg):

(y-0)=-2(x-2) y=-2(x)+(-2)(-2) Right....

OpenStudy (anonymous):

so to get the x in the other side i just add 2x?

OpenStudy (debbieg):

So now your equation is: y=-2x+4 Now add 2x to both sides.

OpenStudy (debbieg):

Exactly! and arrange the left side so that the x term is first, so you'll have Ax+By=C.

OpenStudy (anonymous):

i got it thanks for the hlp debbieg i really apreciate it :P

OpenStudy (debbieg):

This is called Standard Form. Here, A B and C already worked out to be integers. But if you had any fractions, you would just multiply both sides by the LCD and that would take care of that. If you had an A<0, you would multiply both sides by (-1) and that would take care of that. Other than that, this is the general idea for putting any linear equation into standard form. :) and you're welcome, happy to help!